Ricky drove from Town A to Town B in 3 hours. His return trip from Town B to Town A took 5 hours because he drove 15 miles per hour slower on the return trip. How fast did Ricky drive on the trip from Town A to Town B?
Please explain step by step how you get the answer, I am having a really hard time grasping distance problems. I know d=rt. I was thinking something like 3x and 5x-15, but I'm not sure how to set it up. PLEASE help me. And any tips you have for distance problems would be greatly appreciated. Thank you. :)

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

already answered

Answer by stanbon(75887)

Ricky drove from Town A to Town B in 3 hours. His return trip from Town B to Town A took 5 hours because he drove 15 miles per hour slower on the return trip. How fast did Ricky drive on the trip from Town A to Town B?
-------
A to B DATA:
time = 3 hrs ; rate = x mph ; distance = r*t = 3x miles
-----
B to A DATA:
time = 5 hrs ; rate = x-15 mph ; distance = r*t = 5(x-15) = 5x-75 miles
--------
Equation:
distance = distance
-------------
3x = 5x-75
2x = 75
x = 37.5 mph (rate from A to B)
===================================
Cheers,
Stan H.
================